Find the different laurent series in the corresponding domains: 1 (2-1) (2-2) ,0< 12-11 <1;1<12-21 <.
Answer #1

90S 0 Page Date Gesien = (z-1) (-2) 2-1) - (2-2) fZ=12 (2-2) (2-2) (2) I 0 <12-11 Kitton R & Z School 2 thus, we lan use geovThis for domain Of 12-115 / laurant series given by 60 S (2-1) (2-1) (2-2) Z-1 nao 1 +(2-1)*+(3=17*...] < (2-1h < (z-21 <a 의Page Date D Applying geomet cas os/ 2-21 S mit 25 01 Bina Št Z-2 Z-2 nco 2.11. 2 7 (122 Z-2 22-2 L 2-2 2-2 06 Σ D=4Z2 (4) (15